OpenStudy (anonymous):

Solve 5x + 14 - 2x = 9 - (4x + 2)

OpenStudy (anonymous):

3x+14=9-4x-2 3x+14=-4x+7 7x=-7 x=-1

OpenStudy (anonymous):

dpalnc will you help with my question!!

OpenStudy (anonymous):

hello?